A home has a triangular backyard. The second angle of the triangle is 4degrees more than the first angle. The third angle is 8degrees more than two times the first angle. Find the angles of the triangular yard.
The measure of the first angle is

Answer by Cromlix(4381)

Hi,
Make first angle = x
Second angle = x + 4
Third angle = 8 + 2x
Total degrees in a triangle = 180 degrees
Total up
x + x + 4 + 8 + 2x = 180
4x + 12 = 180
4x = 180 -12
4x = 168
x = 168/4
x = 42
Measure of first angle is 42 degrees
Measure of the second angle is 46 degrees
Measure of third angle is 92 degrees
Hope this helps :-)